# Dielectronic recombination studies of ions relevant to kilonovae and non-LTE plasma

### v1

@@ Abstract
This study presents calculations of rate coefficients, resonance strengths, and cross sections for the dielectronic recombination (DR) of $\text{Y}^{+}$, $\text{Sr}^{+}$, $\text{Te}^{2+}$, and $\text{Ce}^{2+}$—low-charge ions relevant to kilonovae and non-local thermodynamic equilibrium (non-LTE) plasmas. Using relativistic atomic structure methods, we computed DR rate coefficients under conditions typical of these environments. Our results highlight the critical role of low-lying DR resonances in shaping rate coefficients at kilonova temperatures ($\sim 10^{4}$ K) and regulating charge-state distributions. Pronounced near-threshold DR resonances significantly influence the evolving ionization states and opacity of neutron star merger ejecta. Comparisons with previous studies emphasize the necessity of including high-$n$ Rydberg states for accurate DR rate coefficients, especially for complex heavy ions with dense energy levels. Discrepancies with existing datasets underscore the need for refined computational techniques to minimize uncertainties. These results provide essential input for interpreting spectroscopic observations of neutron star mergers, including James Webb Space Telescope data. We also put forward suitable candidates for experimental studies, recognizing the challenges involved in such measurements. The data presented here have potential to refine models of heavy-element nucleosynthesis, enhance plasma simulation accuracy, and improve non-LTE plasma modeling in astrophysical and laboratory settings.
@@ I Introduction
Dielectronic recombination (DR), i.e. the capture of a free electron via a simultaneous excitation of a bound electron followed by radiative stabilization is an atomic process that affects the ionization balance, radiative cooling, and energy distribution in both astrophysical and laboratory plasmas, particularly under non-local thermodynamic equilibrium (non-LTE) conditions [1, 2, 3, 4, 5, 6, 7, 8, 9]. In non-LTE conditions, where ionization states are not governed by local thermodynamic equilibrium, the interplay between ionization and recombination processes like DR is critical for determining plasma dynamics. This importance extends across a wide range of astrophysical and laboratory plasmas, particularly in systems characterized by diverse ionization states. An accurate treatment of DR is crucial for improving plasma models and interpreting observational data.
@@
In laboratory plasmas, DR significantly influences energy loss mechanisms, contributing to radiative cooling and the stability of high-temperature plasmas, such as those in fusion research. The sensitivity of high Rydberg states to weak electric fields enhances DR rates even at low plasma densities, making it relevant to environments like the solar corona and magnetic confinement fusion devices [4]. By facilitating energy release through photon emission, DR aids in regulating plasma temperatures and energy distribution, critical factors for both experimental and theoretical studies of plasma behavior.
@@
Furthermore, DR is integral to the development of collisional-radiative models, which describe the interactions between charged particles and photons in non-LTE plasmas [10, 11]. These models are essential for predicting the response of plasmas to varying conditions, including the effects of external fields that can modify ionization potentials and recombination rates. By incorporating DR, these models provide insights into phenomena such as Auger electron heating, hot electron instabilities, and ionization potential depression, which are vital for understanding plasma diagnostics and advancing non-equilibrium plasma theory [12, 13, 14].
@@
While DR has been extensively studied in various contexts, including atomic structure and isotope shifts [15, 16, 17], its impact on kilonova-relevant ions remains an area of active investigation. Astrophysical plasmas, particularly those found in kilonovae, present extreme environments where non-LTE conditions govern the ionization structure and radiative processes. Kilonovae are bright, short-lived transients that occur following the merger of neutron stars, producing heavy elements through rapid neutron-capture (r-process) nucleosynthesis. These explosions eject high-energy material into space, evolving under extreme temperatures and densities, making them ideal laboratories for studying non-LTE plasmas. Within such environments, DR plays a dominant role in electron-ion recombination, significantly influencing ionization balance and spectral evolution. Due to its involvement with high Rydberg states, DR is highly sensitive to density effects, which govern redistributive collisions before radiative stabilization occurs. In kilonovae, which serve as key sites for rapid neutron-capture (r-process) nucleosynthesis at temperatures of approximately $10^{4}$ K, DR often competes with or even surpasses radiative recombination (RR) in shaping the ionization structure of the ejecta. Under such extreme conditions, DR becomes the predominant recombination mechanism for many heavy ions, surpassing RR [18]. This highlights the critical role of DR in shaping the ionization structure of kilonova ejecta, influencing both transient emission and the synthesis of heavy elements, mostly elements that are much heavier than Fe [19, 20, 21].
@@
As the kilonova evolves from the diffusion phase to the nebular phase, recombination alters ionization fractions, leading to changes in spectral line intensities and the emergence or suppression of features in the optical and near-infrared regions. The rate coefficients for DR in r-process elements exhibit strong temperature dependence, particularly for singly to doubly ionized species at $T\sim 10^{4}$ K. Even slight variations in ejecta conditions can lead to substantial differences in ionization balance, directly affecting kilonova spectra. In contrast, RR typically dominates in lower-temperature environments, such as those encountered during the nebular phase of supernovae. Therefore, accurate temperature-dependent DR rates are essential for reliable spectral modeling and the interpretation of transient astrophysical events.
@@
The importance of DR in kilonovae also stems from the complex electron shell structures of heavy elements, which result in dense energy levels and numerous autoionizing states. These states enhance the efficiency of DR. The photon emission in DR during radiative stabilization, known as dielectronic satellite emission, serves as a crucial diagnostic tool for interpreting kilonova spectra and constraining plasma parameters such as temperature and electron density. As a result, DR-driven emission plays a vital role in X-ray spectroscopy and plasma diagnostics, offering valuable insights into the physical conditions of astrophysical plasmas [4, 22].
@@
Low-charged heavy ions such as $\text{Y}^{+},\text{Sr}^{+},\text{Te}^{2+},\text{Ce}^{2+}$ [23, 24, 25] have strong relevance to kilonovae. Observations by the James Webb Space Telescope have detected doubly charged tellurium in GRB 230307A, while singly charged strontium was detected in kilonova AT2017gfo, highlighting the importance of accurate recombination data for understanding low-charge ions in astrophysical plasmas [24, 23]. Heavy ions with excited states near the ground state (e.g., Y+, Sr+, and Ce2+) exhibit complex DR behavior even at low densities, leading to discrepancies in ionization balance calculations. Additionally, different elements contribute uniquely to kilonova opacity and spectral features; Sr and Y produce strong optical and near-infrared transitions, making them detectable in kilonova spectra, while Te and Ce, with their intricate electronic structures, significantly affect mid-infrared opacities. The recombination rates of these elements govern their ionization fractions over time, directly shaping the spectral evolution of kilonovae. Thus, incorporating precise recombination data is essential for improving kilonova models and accurately interpreting their spectral signatures.
@@
Accurate data on DR rates, oscillator strengths, and cross-sections for electron impact are crucial for quantitative analyses of astrophysical spectra, particularly in kilonova ejecta and non-LTE plasmas. However, the scarcity of DR data for key r-process elements such as Y, Sr, Te, and Ce introduces uncertainties in modeling ionization balance and spectral evolution. The dense energy level structures of heavy elements further complicate direct comparisons of recombination rates, as these rates can vary significantly across different ions and even among transitions within the same species. Due to the experimental challenges in measuring DR rates for low-charged heavy elements, most astrophysical models rely on extrapolations from lighter elements [26] or simplified approximations, which can introduce significant inaccuracies in simulations. In particular, models of kilonovae and non-LTE plasmas often depend on crude estimates for low-charged heavy ions, where discrepancies from actual values can reach an order of magnitude, directly impacting astrophysical predictions. It is also important to note that the theoretical study of low-charge heavy ions is rather challenging. While numerous studies have focused on highly charged heavy ions [27, 28, 29, 30, 31, 32, 33, 34, 35], research on low-charge heavy ions remains very limited. To address this gap, the present work employs state-of-the-art atomic structure calculations to provide reliable DR rate coefficients for such ions. These theoretical benchmarks are essential for improving the accuracy of simulation models, refining ionization balance calculations, and enhancing the interpretation of observational data from neutron star mergers and other transient astrophysical events.
@@
To the best of our knowledge, this work represents one of the first computations for the present targets. No experimental measurements have been found in the literature, and the only theoretical calculations available are from a recent study by Banerjee et al. [36], which covers only a few of the ions studied in this work. Another key objective is to provide reference data for experimentalists to facilitate state-of-the-art research in this field. Recent experimental advancements have enabled the study of low-charge heavy ions at the cryogenic storage ring of the Max Planck Institute for Nuclear Physics in Heidelberg [37]. Notably, successful experiments have been conducted on Xe3 ions, marking one of the highest mass-to-charge ratio DR experiments ever performed. However, such experimental efforts face numerous challenges, including the difficulty of generating such ion sources in the laboratory, managing high background noise, and detecting weak DR signals, which are often lost in the background [37]. Given the current limitations of laboratory experiments, theoretical calculations such as those presented here are required. They provide critical insights and benchmarks not only for experimentalists striving to overcome these challenges but also for astrophysicists aiming to model and interpret complex astrophysical phenomena with greater accuracy.
@@ II Theoretical Calculations
For a dielectronic recombination channel, i.e. for a two-step transition $i\to d\to f$, the cross section is expressed as a function of the electron kinetic energy $E$ in the independent resonances approximation as (see, e.g. [38, 39, 40, 41, 32])
@@
derived by summing across all possible autoionization channels and averaging over the Maxwellian distribution of electron energies [42]. In this expression, $k$ denotes the Boltzmann constant, $h$ is the Planck constant that we write explicitly here, $T$ represents the electron temperature, and $E$ is the resonance energy.
@@ III Results and Discussion
In this study, the relativistic configuration interaction method with independent-particle basis wave functions was utilized to calculate the energy levels, radiative rates, autoionization rates, and DR cross sections and rates as implemented in the Flexible Atomic Code (FAC) [43, 44, 45]. The relativistic distorted-wave approximation was employed to describe the continuum states. For the calculation of wave functions and energy levels corresponding to the initial states, intermediate doubly excited states, and radiative final states, contributions from electron correlations, quantum electrodynamics (QED) effects, and Breit interactions were systematically accounted for, ensuring accurate determination of energy levels and wave functions.
@@
Figure 1 illustrates the total DR cross sections for the recombination of Y+ into Y. The resolution width is set to 100 meV, modeled as a Lorentzian profile convoluted with a Gaussian. The calculations include Rydberg states up to $n=6,\ldots,35$. A strong DR resonance is observed near the threshold, accompanied by relatively weaker resonance peaks at approximately 0.3 eV and 0.6 eV. These peaks are attributed to the $4d5s~{}^{3}D_{J}$ states (where $J=1,2,3$). Additionally, a small resonance around 2 eV is associated with the $5s5p~{}^{3}P_{J}$ states (where $J=0,1,2$).
The DR resonance strength for Y+ is approximately an order of magnitude higher than that of previous ions studied by our group [45, 37]. The presence of strong resonances at low energies suggests that DR is dominated by optically allowed core excitations, and it is a dominant recombination mechanism for Y+ under various plasma conditions, particularly in astrophysical environments. Given the exceptionally strong resonance near the threshold, Y+ is an excellent candidate for experimental validation, as background effects are expected to be minimal, enhancing detection feasibility. However, experimental challenges remain, particularly in distinguishing the DR signal from radiative recombination, which is typically prominent at near-threshold energies. Effectively separating these contributions will be crucial for future measurements. The findings of this work could facilitate both observational and experimental studies of Y+ in the near future.
@@
The total DR spectrum for Sr+ recombining to neutral Sr is presented in Figure 2, with a resolution width set to 100 meV. The DR cross section exhibits characteristic resonance peaks, arising from the capture of free electrons into autoionizing Rydberg states via core excitations ($5s\to 5p,4d$). A strong DR resonance feature is observed near zero energy; however, its exact origin remains unidentified. There is a strong possibility that it results from a theoretical artifact associated with the Rydberg state $n=6$, as observed in our calculations, since no such excited states have been reported in the literature at such low energies. Another prominent resonance appears at approximately 2.3 eV, primarily originating from the $4p^{6}4d~{}^{2}D_{5/2}$ and ${}^{2}D_{3/2}$ states, with the ${}^{2}D_{5/2}$ state being the dominant contributor. Additionally, a strong DR resonance is observed around 2.7 eV, attributed to the $4p^{6}5p~{}^{2}P_{1/2}$ and ${}^{2}P_{3/2}$ states. These states are also responsible for the knee-like resonances observed around 3 eV.
Among the ions studied in this work, the DR spectrum of Sr+ exhibits the weakest signal strength, possibly presenting significant challenges for experimental observation with currently available setups. Its DR signal strength is approximately one to two orders of magnitude smaller than that of the other ions studied in this work as well as in previous works[45, 37]. To detect this spectrum experimentally, it would be necessary to significantly reduce background noise.
@@
Figure 3 presents the total DR cross section for Te2+ recombining into Te+, calculated with a resolution width of 100 meV. In the DR spectrum, a small resonance is observed at low energies, below 0.5 eV, attributed to the $5s^{2}5p^{2}~{}^{3}P_{J}$ states (where $J=0,1,2$). At approximately 14.5 eV, two prominent resonance structures appear, corresponding to the $5p5d~{}^{3}D_{J}$ (where $J=1,2,3$) and ${}^{3}P_{J}$ (where $J=0,1,2$) states, with the ${}^{3}D_{1}$ and ${}^{3}P_{1}$ states being the dominant contributors. Additionally, several significant resonance peaks are observed on either side of this strong resonance feature, arising from the $5s5p^{3}$ states.
@@
Figure 4 presents the total dielectronic recombination (DR) cross sections for the recombination of Ce2+ into Ce+, calculated with a resolution width of 100 meV. A strong DR resonance is observed near the threshold, primarily due to the convolution of the $4f^{2}$ and $4f~{}5d$ states. At energies close to 0 eV, the $4f^{2}~{}^{3}H_{5}$ state is the dominant contributor, whereas toward the tail end of the peak, the $4f~{}5d~{}^{1}G_{4}$ state becomes the most significant. Although the spectrum appears to decline beyond 0.2 eV, the cross-section magnitude remains on the order of $10^{7}$ barns, emphasizing the strength of this prominent resonance feature. This suggests that Ce2+ could serve as a strong candidate for experimental study near the threshold energy, similar to Y+. Furthermore, the exceptionally high DR signal might indicate a considerable probability of its detection in a kilonova.
The findings of Gu [43] suggest that the energy values computed using the FAC may have uncertainties due to factors such as incomplete treatment of correlations or limitations in the atomic model. Furthermore, according to the FAC manual authored by Gu [46], in the case of near-neutral ions (as in the present case), the uncertainties associated with radiative decay rates and autoionization rates can reach 20% or, in certain instances, exceed this value. Additionally, limiting the number of states $n$ introduces an uncertainty of approximately 10%. Therefore, the overall minimum uncertainty in the present calculation of DR cross section and strengths is estimated to be around 30%. Given the substantial amount of data generated in this work, all data are made available as supplementary material.

### v2

@@ Abstract
This study presents calculations of rate coefficients, resonance strengths, and cross sections for the dielectronic recombination (DR) of $\text{Y}^{+}$, $\text{Sr}^{+}$, $\text{Te}^{2+}$, and $\text{Ce}^{2+}$—low-charge ions relevant to kilonovae and non-local thermodynamic equilibrium (non-LTE) plasmas. Using relativistic atomic structure methods, we computed DR rate coefficients under conditions typical of these environments. These DR rate coefficients and cross sections were calculated using the Flexible Atomic Code (FAC). The DR resonance features were identified by comparing theoretical resonance energies, estimated as the difference between NIST (National Institute of Standards and Technology) excitation energies and Dirac binding energies, with dominant autoionizing states confirmed through analysis of autoionization rates. Our results highlight the critical role of low-lying DR resonances in shaping rate coefficients at kilonova temperatures ($\sim 10^{4}$ K) and regulating charge-state distributions. Pronounced near-threshold DR resonances significantly influence the evolving ionization states and opacity of neutron star merger ejecta. Comparisons with previous studies emphasize the necessity of including high-$n$ Rydberg states for accurate DR rate coefficients, especially for complex heavy ions with dense energy levels. Discrepancies with existing datasets underscore the need for refined computational techniques to minimize uncertainties. These results provide essential input for interpreting spectroscopic observations of neutron star mergers, including James Webb Space Telescope data. We also put forward suitable candidates for experimental studies, recognizing the challenges involved in such measurements. The data presented here have the potential to refine models of heavy-element nucleosynthesis, enhance plasma simulation accuracy, and improve non-LTE plasma modeling in astrophysical and laboratory settings.
@@ I Introduction
Dielectronic recombination (DR) affects the ionization balance, radiative cooling, and energy distribution in plasmas, particularly under non-local thermodynamic equilibrium (non-LTE) conditions (1, 2, 3, 4, 5, 6, 7, 8, 9). In such non-LTE environments, where ionization states are not governed by local thermodynamic equilibrium, the interplay between ionization and recombination processes like DR is critical for determining plasma dynamics across both astrophysical and laboratory systems with diverse ionization states. An accurate treatment of DR is crucial for improving plasma models and interpreting observational data.
@@
In laboratory plasmas, DR significantly influences energy loss mechanisms, and enhances plasma stability, particularly in fusion research. The sensitivity of high Rydberg states to weak electric fields enhances DR rates at low plasma densities, making it relevant to environments like the solar corona and magnetic confinement fusion devices (4). By emitting photons, DR helps regulate plasma temperature, a key factor for both experimental investigations and theoretical models.
@@
Furthermore, DR is essential to collisional-radiative models, which describe the interactions between charged particles and photons in non-LTE plasmas (10, 11). These models predict plasma responses to varying conditions, including external field effects on ionization potentials and recombination rates. Incorporating DR in these models provides insight into Auger electron heating, hot electron instabilities, and ionization potential depression, which are key to plasma diagnostics and non-equilibrium plasma theory (12, 13, 14).
@@
DR studies on kilonova-relevant ions remain an area of active investigation. Astrophysical plasmas in kilonovae present extreme, non-LTE environments where ionization structure and radiative processes are strongly influenced. DR plays a dominant role in electron-ion recombination, significantly influencing spectral evolution. Involving high Rydberg states, DR is highly sensitive to density effects, which govern redistributive collisions before radiative stabilization occurs. In kilonovae, which serve as key sites for rapid neutron-capture (r-process) nucleosynthesis at temperatures of approximately $10^{4}$ K, DR often competes with or even surpasses radiative recombination (RR) in shaping the ionization structure of the ejecta. Under such extreme conditions, DR becomes the predominant recombination mechanism for many heavy ions, surpassing RR (15). This underscores DR’s critical role in determining kilonova ejecta’s ionization structure, transient emission, and the synthesis of elements much heavier than Fe (16, 17, 18).
@@
As a kilonova evolves from the diffusion to the nebular phase, recombination alters ionization fractions, changing spectral line intensities and causing the emergence or suppression of features in the optical and near-infrared regions. DR rate coefficients in r-process elements exhibit strong temperature dependence, particularly for singly to doubly ionized species at $T\sim 10^{4}$ K. Even slight variations in ejecta conditions can significantly affect ionization balance and, consequently, kilonova spectra. In contrast, RR tends to dominate in lower-temperature environments, such as those in the nebular phase of supernovae. Thus, accurate temperature-dependent DR rates are essential for reliable spectral modeling and interpreting transient astrophysical events.
@@
The importance of DR in kilonovae also stems from the complex electron shell structures of heavy elements, which have dense energy levels and numerous autoionizing states, enhancing DR efficiency. The photon emission in DR during radiative stabilization, serves as a crucial diagnostic tool for interpreting kilonova spectra and constraining plasma parameters such as temperature and electron density. As a result, DR-driven emission plays a vital role in X-ray spectroscopy and plasma diagnostics, offering valuable insights into the physical conditions of astrophysical plasmas (4, 19).
@@
Low-charged heavy ions such as $\text{Y}^{+},\text{Sr}^{+},\text{Te}^{2+},\text{Ce}^{2+}$ (20, 21, 22) are highly relevant to kilonovae. The James Webb Space Telescope has detected Te2+ in GRB 230307A, and Sr+ was detected in kilonova AT2017gfo, underscoring the need for accurate recombination data to understand low-charge ions in astrophysical plasmas (21, 20). Heavy ions with low-lying excited states (e.g., Y+, Sr+, and Ce2+) exhibit complex DR behavior even at low densities, causing discrepancies in ionization balance calculations. Moreover, different elements affect kilonova opacity and spectra uniquely: Sr and Y produce strong optical and near-infrared transitions, making them observable, while Te and Ce, with their intricate electronic structures, significantly affect mid-infrared opacities. Their recombination rates determine ionization fractions over time, shaping kilonova spectral evolution. Thus, precise recombination data are essential for improving kilonova models and interpreting spectral signatures accurately.
@@
Accurate data on DR rates, oscillator strengths, and electron-impact cross-sections are essential for quantitative analyses of astrophysical spectra, particularly in kilonova ejecta and non-LTE plasmas. However, the scarcity of DR data for key r-process elements like Y, Sr, Te, and Ce introduces uncertainties in modeling ionization balance and spectral evolution. The dense energy level structures of heavy elements further complicate direct comparisons, as recombination rates can vary significantly across ions and even between transitions within the same species. Due to experimental challenges in measuring DR rates for low-charged heavy elements, models often rely on extrapolations from lighter elements (23) or simplified approximations, leading to substantial inaccuracies. In particular, kilonovae and non-LTE plasma models often depend on crude estimates for low-charged heavy ions, where discrepancies can reach an order of magnitude, directly impacting astrophysical predictions. Theoretical studies of these ions are also difficult: while highly charged heavy ions have been widely studied (24, 25, 26, 27, 28, 29, 30, 31, 32), research on low-charge species remains limited. To address this gap, the present work uses state-of-the-art atomic structure calculations to provide reliable DR rate coefficients for such ions. These theoretical benchmarks are essential for improving simulation accuracy, refining ionization balance calculations, and enhancing the interpretation of observational data from neutron star mergers and other transient astrophysical events.
@@
To the best of our knowledge, this work presents one of the first computations for the current targets. No experimental measurements exist in the literature, and the only available theoretical data come from a recent study by Banerjee et al. [33], which includes only a few of the ions considered here. Another key objective is to provide reference data for experimentalists to facilitate state-of-the-art research in this field. Recent experimental advancements have enabled the study of low-charge heavy ions at the cryogenic storage ring of the Max Planck Institute for Nuclear Physics in Heidelberg (34). Notably, successful DR measurements on Xe3+ represent the highest mass-to-charge ratio experiments conducted to date. However, these experiments face major challenges, including difficulties in ion source generation, high background noise, and weak DR signals that are often obscured (34). Given these experimental limitations, theoretical calculations such as those presented here are required. They provide critical insights and benchmarks not only for experimentalists striving to overcome these challenges but also for astrophysicists modeling and interpreting complex astrophysical phenomena with greater accuracy.
@@ II Theoretical calculations
For a dielectronic recombination channel, i.e. for a two-step transition $i\to d\to f$, the cross section is expressed as a function of the electron kinetic energy $E$ in the independent resonances approximation as (see, e.g. 35, 36, 37, 38, 29)
@@
derived by summing across all possible autoionization channels and averaging over the Maxwellian distribution of electron energies (39). In this expression, $k$ denotes the Boltzmann constant, $h$ is the Planck constant that we write explicitly here, $T$ represents the electron temperature, and $E$ is the resonance energy.
@@ III Results and discussion
In this study, the relativistic configuration interaction method with independent-particle basis wave functions was utilized to calculate the energy levels, radiative rates, autoionization rates, and DR cross sections and rates as implemented in the Flexible Atomic Code (FAC) (40, 41, 42). The relativistic distorted-wave approximation was employed to describe the continuum states. For the calculation of wave functions and energy levels corresponding to the initial states, intermediate doubly excited states, and radiative final states, contributions from electron correlations, quantum electrodynamics (QED) effects, and Breit interactions were systematically accounted for, ensuring accurate determination of energy levels and wave functions.
The DR resonance features were identified by comparing the theoretically calculated resonance energies with known excitation energies of the core ion, as listed in the NIST Atomic Spectra Database. The DR resonance energy, $E_{\text{DR}}$, is approximately given by the difference between the excitation energy of the bound core electron, $E_{\text{ex}}$, and the binding energy of the captured free electron, $E_{\text{B}}$; that is, EDR = Eex - EB. Here, $E_{\text{ex}}$ values were taken from NIST, while $E_{\text{B}}$ was estimated using Dirac binding energies (43). This provides an approximate position for the expected DR resonance features. To further verify these identifications, we analyzed the autoionization rates of the intermediate states. Since higher autoionization rates typically correspond to stronger DR resonance features, the dominant contributors identified through rate analysis were found to be consistent with the states inferred from energy comparisons.
@@
Figure 1 illustrates the total DR cross sections for the recombination of Y+ into Y. The resolution width is set to 100 meV, modeled as a Lorentzian profile convoluted with a Gaussian. The calculations include Rydberg states up to $n=6,\ldots,35$. A strong DR resonance is observed near the threshold, accompanied by relatively weaker resonance peaks at approximately 0.3 eV and 0.6 eV. These peaks are attributed to the $4d5s~{}^{3}D_{J}$ states (where $J=1,2,3$). Additionally, a small resonance around 2 eV is associated with the $5s5p~{}^{3}P^{o}_{J}$ states (where $J=0,1,2$).
The DR resonance strength for Y+ is approximately an order of magnitude higher than that of previous ions studied by our group (42, 34). The presence of strong resonances at low energies suggests that DR is dominated by optically allowed core excitations, and it is a dominant recombination mechanism for Y+ under various plasma conditions, particularly in astrophysical environments. Given the exceptionally strong resonance near the threshold, Y+ is an excellent candidate for experimental validation, as background effects are expected to be minimal, enhancing detection feasibility. However, experimental challenges remain, particularly in distinguishing the DR signal from radiative recombination, which is typically prominent at near-threshold energies. Effectively separating these contributions will be crucial for future measurements. The findings of this work could facilitate both observational and experimental studies of Y+ in the near future.
@@
The total DR spectrum for Sr+ recombining to neutral Sr is presented in Figure 2, with a resolution width set to 100 meV. The DR cross section exhibits characteristic resonance peaks, arising from the capture of free electrons into autoionizing Rydberg states via core excitations ($5s\to 5p,4d$). A strong DR resonance feature is observed near zero energy; however, its exact origin remains unidentified. There is a strong possibility that it results from a theoretical artifact associated with the Rydberg state $n=6$, as observed in our calculations, since no such excited states have been reported in the literature at such low energies. Another prominent resonance appears at approximately 2.3 eV, primarily originating from the $4p^{6}4d~{}^{2}D_{5/2}$ and ${}^{2}D_{3/2}$ states, with the ${}^{2}D_{5/2}$ state being the dominant contributor. Additionally, a strong DR resonance is observed around 2.7 eV, attributed to the $4p^{6}5p~{}^{2}P^{o}_{1/2}$ and ${}^{2}P^{o}_{3/2}$ states. These states are also responsible for the knee-like resonances observed around 3 eV.
Among the ions studied in this work, the DR spectrum of Sr+ exhibits the weakest signal strength, possibly presenting significant challenges for experimental observation with currently available setups. Its DR signal strength is approximately one to two orders of magnitude smaller than that of the other ions studied in this work as well as in previous works (42, 34). To detect this spectrum experimentally, it would be necessary to significantly reduce background noise.
@@
Figure 3 presents the total DR cross section for Te2+ recombining into Te+, calculated with a resolution width of 100 meV. In the DR spectrum, a small resonance is observed at low energies, below 0.5 eV, attributed to the $5s^{2}5p^{2}~{}^{3}P_{J}$ states (where $J=0,1,2$). At approximately 14.5 eV, two prominent resonance structures appear, corresponding to the $5p5d~{}^{3}D^{o}_{J}$ (where $J=1,2,3$) and ${}^{3}P^{o}_{J}$ (where $J=0,1,2$) states, with the ${}^{3}D^{o}_{1}$ and ${}^{3}P^{o}_{1}$ states being the dominant contributors. Additionally, several significant resonance peaks are observed on either side of this strong resonance feature, arising from the $5s5p^{3}$ states.
@@
Figure 4 presents the total dielectronic recombination (DR) cross sections for the recombination of Ce2+ into Ce+, calculated with a resolution width of 100 meV. A strong DR resonance is observed near the threshold, primarily due to the convolution of the $4f^{2}$ and $4f~{}5d$ states. At energies close to 0 eV, the $4f^{2}~{}^{3}H_{5}$ state is the dominant contributor, whereas toward the tail end of the peak, the $4f5d^{1}G^{o}_{4}$ state becomes the most significant. As seen in Table 1, although the excitation energy of the $4f5d^{1}G^{o}_{4}$ state appears lower than that of the $4f^{2}~{}^{3}H_{5}$ state in the FAC calculations, the inclusion of the Rydberg electron introduces additional configuration mixing. As a result, the resonance energy associated with the $4f5d^{1}G^{o}_{4}$ state becomes higher than that of the $4f^{2}~{}^{3}H_{5}$ state. The contribution of the $4f5d^{1}G^{o}_{4}$ state at the tail end of the resonance feature is further supported by the analysis of the autoionization rates. Although the spectrum appears to decline beyond 0.2 eV, the cross-section magnitude remains on the order of $10^{7}$ barns, emphasizing the strength of this prominent resonance feature. This suggests that Ce2+ could serve as a strong candidate for experimental study near the threshold energy, similar to Y+. Furthermore, the exceptionally high DR signal might indicate a considerable probability of its detection in a kilonova.
Table 1: Comparison of energy levels (in eV) from present FAC calculations and NIST (44) for selected ions.
Ion State FAC NIST
Y+ 4d5s2 2D3/2 (Y) 0 0
5s2 1S0 4.728 6.217
4d5s 3D1 4.978 6.321
4d5s 3D2 4.994 6.347
4d5s 3D3 5.021 6.397
[…]
Sr+ 5s2 1S0 (Sr) 0 0
4p65s 2S1/2 4.603 5.695
4p64d 2D3/2 6.878 7.500
4p64d 2D5/2 6.879 7.534
[…]
5s25p2 3P0 17.798 18.600
5s25p2 3P1 18.274 19.189
5s25p2 3P2 18.740 19.612
[…]
5s25p5d ${}^{\circ}_{3}$ 32.466 33.589
[…]
4f2 3H4 12.388 10.956
4f2 3H5 12.498 11.145
4f2 3H6 12.642 11.344
[…]
Table 1 presents a comparison between the energy levels calculated in this work using the FAC and the values available from the NIST Atomic Spectra Database (44). As noted by [40], energy levels computed using FAC can exhibit uncertainties due to incomplete treatment of correlations or limitations in the atomic model or lack of configuration mixing. From Table 1, it is evident that the calculated energy levels deviate from the NIST values by approximately 1 to 1.5 eV. Such differences are expected for near-neutral heavy ions, as the estimated uncertainty in energy levels calculated using FAC can be of the order of a few electronvolts. A recent study by [45], which also investigates low-charged lanthanides relevant to kilonovae, reports similar discrepancies of a few eVs when comparing theoretical values with NIST data. This supports the observation that such deviations are typical when modeling complex atomic structures. In addition to energy uncertainties, there are inherent uncertainties in the computed rates and cross sections. According to the FAC manual authored by [46], in the case of near-neutral ions (as in the present case), the uncertainties associated with radiative decay rates and autoionization rates can reach 20% or, in certain instances, exceed this value. Additionally, limiting the number of states $n$ introduces an uncertainty of approximately 10%. Therefore, the overall uncertainty in the present calculation of DR cross section and strengths is estimated to be around 30%. Given the substantial amount of data generated in this work, all data are made available as supplementary material.